OpenStudy (anonymous):

c + 6 < –20 help sole the inequality

OpenStudy (anonymous):

c + 6 < -20

OpenStudy (anonymous):

c < -26

OpenStudy (anonymous):

c<-26

OpenStudy (anonymous):

\[\frac{ 9 }{ 8 } \ge -5\] help me please

OpenStudy (anonymous):

C<-26

OpenStudy (anonymous):

C< -40/9

OpenStudy (anonymous):

\[\frac{ p }{ ?8}\] sorry

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

\[\frac{ P }{ 8 } \ge -5\]

OpenStudy (anonymous):

p is greater than or equal to -40

OpenStudy (anonymous):

-8x > -64

OpenStudy (anonymous):

x<8

OpenStudy (anonymous):

@Lakeekee_

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

what that website again?

OpenStudy (anonymous):

@Lakeekee_

OpenStudy (anonymous):

www.mathway.com